An elephant lifts a fallen tree by exerting a force of 8.25 x 103 N upward. If she lifts the tree 2.10 m in 1.05 seconds, how much power does she use?
1
Expert's answer
2020-11-20T09:32:29-0500

By definition, the power is:


"P = \dfrac{F\cdot s}{t}"

where "F = 8250N" is the exerted force, "s = 2.1m" is the distance and "t = 1.05s" is a time. Thus, obtain:


"P = \dfrac{8250\cdot 2.1}{1.05} =16500\space W = 16.5\space kW"

Answer. 16500 W or 16.5 kW.